# Robust Concept Erasure in Diffusion Models: A Theoretical Perspective on Security and Robustness

### v1

@@ 2.1 Concept Erasure in Generative Models
Concept erasure has emerged as a critical problem in generative AI. ESD (Erase Stable Diffusion) (Gandikota et al., 2023) first introduced the idea of fine-tuning diffusion models with negative prompts, steering generations away from undesired content. While effective for single concepts, it often induces collateral damage to unrelated content. UCE (Unlearning via Concept Editing) (Gandikota et al., 2024) proposed a closed-form solution for cross-attention weight editing, enabling efficient edits but sacrificing robustness when concepts are entangled.
@@ 2.3 Diffusion-Based Image Editing
Beyond erasure, diffusion models have been widely used for image editing. Prompt-to-Prompt (Hertz et al., 2022) and TF-ICON (Lu et al., 2023) modifies attention maps to adjust generated outputs with fine control. InstructPix2Pix (Brooks et al., 2023) enables natural language editing instructions. Diffusion-based editing excels at user-driven modifications but does not fundamentally remove model-internal knowledge. SCORE differs by permanently altering the model such that the erased concept cannot be generated in any context.
@@ Appendix A Additional Backgrounds
With the advancement of deep learning (Zheng et al., 2024, 2023, 2022, 2025a; Yu et al., 2025c, b; Qiu et al., 2024, 2025d, 2025c; Liu et al., 2025d; Qiu et al., 2025a, b; Wu et al., 2025c, 2024c; Li et al., 2025e; Gao et al., 2025; Hu et al., 2024; Wu et al., 2024b, 2023, a; Luo et al., 2025; Mao et al., 2025; Sun et al., 2025a, b; Niu et al., 2025; Kudrat et al., 2025; Han et al., 2025a, c, d; Zeng et al., 2025; Han et al., 2025b; Feng et al., 2024; Huang et al., 2025a; Xie et al., 2025; Tang et al., 2022, 2025a, 2023; Shan et al., 2021; Fang et al., 2020; Cui et al., 2021; Shan et al., 2022; Hu et al., 2025; Nie et al., 2025; Zhou et al., 2023, 2025, 2024; Shi et al., 2025; Zhao et al., 2025; Li et al., 2025d, c; Chen et al., 2025c; Huang et al., 2025b; Fu et al., 2025; Gong et al., 2021, 2022, 2024b, 2024a; Bi et al., 2024, 2025d, 2025c; Wang et al., 2025; Chen et al., 2025b, a; Rong et al., 2025; Zhang et al., 2023; Zheng et al., 2024, 2023, 2022, 2025a; Yue et al., 2025; Lin et al., 2024a; Huang et al., 2025c, e, d; Lin et al., 2024b; Liu et al., 2024c; Bi et al., 2025a; Tang et al., 2025d; Bi et al., 2025b; Tang et al., 2025b; Liu et al., 2025b, a; Zhang et al., 2025a; Song et al., 2024b; Liu et al., 2025b; Song et al., 2024a; Liu et al., 2025c; Tang et al., 2025c; Liu et al., 2024b; Tang et al., 2024; Song et al., 2024c; Liu et al., 2024a; Lou et al., 2023; Li et al., 2024c, a, b, 2023, 2022; Guo et al., 2023b, a; Liang et al., 2022; Yu et al., 2025a; Shen et al., 2025; Zhang et al., 2025b, c, d; Lin & Losavio, 2025; Jiang et al., 2025; Li et al., 2025b; Lin et al., 2025; Lin & Shih, ; Peng et al., 2024c; Di et al., 2025; Peng et al., 2024a; Feng et al., 2025; Wang et al., 2023b, a; Peng et al., 2021, 2020; Wu et al., 2025a; He et al., 2024c; Yang et al., 2022a, b; Zheng et al., 2025b) and generative models (He et al., 2025c, b, 2023a, d, e, a, 2024a, 2024b, 2023c, 2023b; Xiao et al., 2024; Wu et al., 2025b; Peng et al., 2025a, 2024b, b, 2024d, c), an increasing number of studies have begun to focus on the issue of concept erasure in generative models.

### v2

@@ 2.1 Concept Erasure in Generative Models
Concept erasure (Gao et al., 2025a) has emerged as a critical problem in generative AI. ESD (Erase Stable Diffusion) (Gandikota et al., 2023) first introduced the idea of fine-tuning diffusion models with negative prompts, steering generations away from undesired content. While effective for single concepts, it often induces collateral damage to unrelated content. UCE (Unlearning via Concept Editing) (Gandikota et al., 2024) proposed a closed-form solution for cross-attention weight editing, enabling efficient edits but sacrificing robustness when concepts are entangled.
@@ 2.3 Diffusion-Based Image Editing
Beyond erasure, diffusion models have been widely used for image editing (Lu et al., 2025; Zhou et al., 2025b). Prompt-to-Prompt (Hertz et al., 2022) and TF-ICON (Lu et al., 2023) modifies attention maps to adjust generated outputs with fine control. InstructPix2Pix (Brooks et al., 2023) enables natural language editing instructions. Diffusion-based editing excels at user-driven modifications but does not fundamentally remove model-internal knowledge. SCORE differs by permanently altering the model such that the erased concept cannot be generated in any context.
@@ Appendix A Additional Backgrounds
With the advancement of deep learning (Zheng et al., 2024, 2023, 2022, 2025a; Yu et al., 2025c, b; Qiu et al., 2024, 2025d, 2025c; Liu et al., 2025d; Qiu et al., 2025a, b; Wu et al., 2025c, 2024c; Li et al., 2025e; Gao et al., 2025b; Hu et al., 2024; Wu et al., 2024b, 2023, a; Luo et al., 2025; Mao et al., 2025; Sun et al., 2025a, b; Niu et al., 2025; Kudrat et al., 2025; Han et al., 2025a, c, d; Zeng et al., 2025; Han et al., 2025b; Feng et al., 2024; Huang et al., 2025a; Xie et al., 2025; Tang et al., 2022, 2025a, 2023; Shan et al., 2021; Fang et al., 2020; Cui et al., 2021; Shan et al., 2022; Hu et al., 2025; Nie et al., 2025; Zhou et al., 2023, 2025a, 2024; Shi et al., 2025; Zhao et al., 2025; Li et al., 2025d, c; Chen et al., 2025c; Huang et al., 2025b; Fu et al., 2025; Gong et al., 2021, 2022, 2024b, 2024a; Bi et al., 2024, 2025d, 2025c; Wang et al., 2025; Chen et al., 2025b, a; Rong et al., 2025; Zhang et al., 2023; Zheng et al., 2024, 2023, 2022, 2025a; Yue et al., 2025; Lin et al., 2024a; Huang et al., 2025c, e, d; Lin et al., 2024b; Liu et al., 2024c; Bi et al., 2025a; Tang et al., 2025d; Bi et al., 2025b; Tang et al., 2025b; Liu et al., 2025b, a; Zhang et al., 2025a; Song et al., 2024b; Liu et al., 2025b; Song et al., 2024a; Liu et al., 2025c; Tang et al., 2025c; Liu et al., 2024b; Tang et al., 2024; Song et al., 2024c; Liu et al., 2024a; Lou et al., 2023; Li et al., 2024c, a, b, 2023, 2022; Guo et al., 2023b, a; Liang et al., 2022; Yu et al., 2025a; Shen et al., 2025; Zhang et al., 2025b, c, d; Lin & Losavio, 2025; Jiang et al., 2025; Li et al., 2025b; Lin et al., 2025; Lin & Shih, ; Peng et al., 2024c; Di et al., 2025; Peng et al., 2024a; Feng et al., 2025; Wang et al., 2023b, a; Peng et al., 2021, 2020; Wu et al., 2025a; He et al., 2024c; Yang et al., 2022a, b; Zheng et al., 2025b) and generative models (He et al., 2025c, b, 2023a, d, e, a, 2024a, 2024b, 2023c, 2023b; Xiao et al., 2024; Wu et al., 2025b; Peng et al., 2025a, 2024b, b, 2024d, c; Qu et al., 2025b, a; Chen & Greer, 2025; Lyu et al., 2025), an increasing number of studies have begun to focus on the issue of concept erasure in generative models.